OpenStudy (anonymous):

PLESSAE HELP ME I WILL REALLY NEED IT I HAVE BEEN WAITING SO LONG I WLL GIVE YOU A FAN AND MEDAL IT IS ALGEBRA AND NO WHAT YOU ARE DOING :D

OpenStudy (anonymous):

Winnie rented a room at $250 per week. She also had to pay a deposit of $1200. Part A: Write a function to show the total amount of money f(n) that Winnie has to pay for renting the room for n weeks. (4 points) Part B: What is the total rent that Winnie has to pay for using the room for 12 weeks? (4 points) Part C: If the deposit amount decreases by $100, what would be the total rent for 6 weeks? (2 points)

OpenStudy (anonymous):

@sleepyjess @TheSmartOne

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

each week she pays 250 so for n weeks she pays \(250n\) add that to the 1200 deposit and you have \[f(n)=250n+1200\]

OpenStudy (anonymous):

So that is the function for part A @satellite73

OpenStudy (anonymous):

@satellite73 @TheSmartOne @Nnesha

OpenStudy (anonymous):

what is c @TheSmartOne @Nnesha @anybody

OpenStudy (anonymous):

well what is part c? i understand the rest

TheSmartOne (thesmartone):

for b just plug in 12 for x

OpenStudy (anonymous):

Part A: Total amount of money is 1200 + $250 a week for n weeks that is simply f(n) = 1200 + 250n Part B: For 12 weeks, that is 1200 + 250(12) = 1200 + 3000 = $4,200

OpenStudy (anonymous):

yeah i know that @TheSmartOne help me w. c please

TheSmartOne (thesmartone):

its just 1200-100

OpenStudy (anonymous):

? @TheSmartOne

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @satellite73 each week she pays 250 so for n weeks she pays \(250n\) add that to the 1200 deposit and you have \[f(n)=250n+1200\] \(\color{blue}{\text{End of Quote}}\) then for part c, its 1200-100 what is 1200-100=?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

why is that though explanationn please why are you subtracting

TheSmartOne (thesmartone):

decreases = subtraction "If the deposit amount decreases by $100"

OpenStudy (anonymous):

so part c is 1,000?

TheSmartOne (thesmartone):

nope

TheSmartOne (thesmartone):

What is \(\sf 1200-100=?\)

OpenStudy (anonymous):

@TheSmartOne please help i have tons of more questions and my mom is going to scream at me if id o not finsih in 30 min. :(

OpenStudy (anonymous):

1200-100=1100

TheSmartOne (thesmartone):

correct, not 1000

OpenStudy (anonymous):

ok so that is the answer for part c?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

do you mind helping me with another ?

TheSmartOne (thesmartone):

@hiopenstudy we are not done with Part C. So now we got 1100. So our new equation will be \(\sf y=250n +1100\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i have to do after this? @TheSmartOne

TheSmartOne (thesmartone):

just plug in 6 for n and solve for y :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y = 2600 @TheSmartOne is that the answer?

TheSmartOne (thesmartone):

Correct :)

OpenStudy (anonymous):

@TheSmartOne PLEASE DO NOT MAKE ME WAIT AGAIN MY MOM AS ALREADY SCREAMED AT ME THT I HAVE NOT FINISHED MY WORK :(